23^2=529-1=528/22=24-23=1
Is this right?
x
x^2
x^2 * 7
(x^2*7) -7
((x^2*7) -7) divided by 3
=x
And...
(3²)-1=8
8÷2=4
4-3=1
(x²)-(x-1)
8÷(x-1)
4-x=1

Answer by richwmiller(17219)

I see what you are trying to do but your technique is not accurate.
23^2=529-1
but 23^2 does NOT equal=529-1
23^2=529
Most certainly 23^2=529-1=528/22=24-23=1 can't be upheld
23^2 does NOT equal 1
You need to show the original problem for us to be able to follow your steps.